five times the quare of a number is 980. What is the number?
i used x^2 + 5x -980
a=1 b=5 c=-980

Answer by jim_thompson5910(35256)

"five times the quare of a number is 980"

translates to

5x^2 = 980

Solve for x

5x^2 = 980

x^2 = 980/5

x^2 = 196

x = sqrt(196) or x = -sqrt(196)

x = 14 or x = -14

Therefore, the number is either 14 or -14.

--------------------------------------------------------------------------------------------------------------

If you need more help, you can email me at jim_thompson5910@hotmail.com

To find other ways to contact me, you can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------